# Errors-In-Variables Model Fitting for Partially Unpaired Data Utilizing Mixture Models

### v1

@@ 2.4.1 Example: Fitting a Line and Gaussian Disturbance (Deming Regression)
representing Deming regression for partially unpaired data.
@@
Gaussian disturbances in input and output variables with ${\boldsymbol{\eta}}_{h}\sim\mathcal{N}(\boldsymbol{0},\,\sigma_{{\eta}}^{2}\cdot I_{k\times k})(\boldsymbol{s})\;\forall h=1,..,H$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$, and utilizing the $k$-dimensional affine model $M(\boldsymbol{x};\boldsymbol{\alpha})=\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{n}$ and inserting it, we get by applying Equation (9)
@@ 2.4.3 Example: Linear Interval Data Regression
As final example, we present the case of interval data that are given with $\boldsymbol{x}_{h}:=\frac{1}{2}\left(\underline{\boldsymbol{x}_{h}}+\overline{\boldsymbol{x}_{h}}\right)$ and $\boldsymbol{y}_{l}:=\frac{1}{2}\left(\underline{\boldsymbol{y}_{l}}+\overline{\boldsymbol{y}_{l}}\right)$, and $v_{h,i}:=\frac{1}{2}\left(\overline{x_{h,i}}-\underline{x_{h,i}}\right)$ and $w_{l,j}:=\frac{1}{2}\left(\overline{y_{l,j}}-\underline{y_{l,j}}\right)$, and
@@ 2.5 Extensions
Numerical implementation strategies for the general formula (9) are presented in Supplement D.
@@ 3.1 Demonstration for a Line Fit
See Figure 2 for presentation of example objective function and resulting line fits of scenario $A$ and $D$. The columns correspond to different numbers of unpaired subgroups $R\in\{1,3,12,L\}$. For the completeley unpaired case ($R=1$) only the objective function is presented.
@@
In Supplement E a plane fit is presented as an example for 2D input variables.
@@ 3.2 Nonlinear Model with Anisotropic Observation Errors
is presented with anisotropic Gaussian disturbances $\sigma_{\eta}=0.2$ and $\sigma_{\varepsilon}=0.1$ for $L=H=300$ data points. In this model, nonlinearity holds with respect to $x$ and not with respect to $\boldsymbol{\alpha}$, which is deliberate and not necessary. In Figure 4 the results are presented for a completely paired case ($R=L=H$) (left) and a partially unpaired case (right) with two areas of lost pairing information. The general algorithm was implement according to section D with the simple trapezoidal rule for integration and the Nelder-Mead-optimization in Matlab.
@@ 4 Real Data Study: Life Expectancy
For evaluation purposes, we perform a train-test-split with $172$ training countries and $20$ test countries in order to judge if we can learn from the training countries the life expectancy for the test countries based on the input variables. A major challenge is that the goodness of fit cannot be measured with the classical $R^{2}$ since it is only valid for cases with no errors in the input variables. In consequence, we utilize the extension of $R^{2}$ to the errors-in-variables approach $R_{\delta}^{2}$ [20] for multiple linear regression (see Supplement F for more details).
First, we consider the case of the model fit with perfect pairing ($R=L=H$). Since the proposed algorithm was implemented with Monte Carlo methods as described in Supplement D, it was verified with an explicit solution for the multiple linear regression model [20] and showed equivalent results. The results are presented in Figure 6A as correlation and error plots of predicted and real output values for the train and test data sets. They show that the training allows a high quality prediction for the test countries with the multiple linear regression model for life expectancy.
@@
Second, one major approach in this work is to investigate explanatory power of the model fit if the train data is partially unpaired, i.e. if we only consider data of groups of countries rather than individual countries for model training. To achieve this, we introduce the country grouping along an additional criteria which is not part of the input: GDP per capita (current USD) [NY.GDP.PCAP.CD] with increasing GDP per capita for each group. Please note, although the grouping is performed by GDP, the GDP itself is not part of the predictors and the overall information level for the predictors is reduced by this grouping compared to the completely paired data set. We considered $3$ different groupings with group sizes $L_{r}=H_{r}$ (the last group always consists of residual countries) of $4$ ($R=44$), $8$ ($R=22$) and $16$ ($R=11$). The results of the model fit based on this country groups are presented in table 1. The $R_{\delta}^{2}$ for train and test data is presented which show overall very high goodness-of-fit values but with decreasing values for increasing group sizes, as one might expect. Please note, the $R_{\delta}^{2}$ for testing contains a rather strong noise component since it is calculated by only $20$ data points in the test set. In addition, in Figure 6B (for group size $16$) it is demonstrated that the fit results still align well for training and testing even for $11$ subgroups. It is noted, that the grouping is part of the training to get $\boldsymbol{\alpha}$, and not part of these scatter plots. This indicates that the model fit and the prediction for the test countries work quite well based only on the country groups, confirming that the pairing information can be reduced and this still provides a valuable model fit.
@@ 5 Discussion and Conclusion
Specifically, we regard the framework presented as a generalization of total least squares approaches [17] as a more general errors-in-variables treatment, since we can reproduce classical results of total least squares by utilizing Gaussian errors and a line model. On the other hand, we are not restricted to these assumptions, e.g. allowing for interval data using uniform distributions. We want to emphasis, that the only initial random variables of this paper are the uncertainty random variables $\boldsymbol{\eta}$ and $\boldsymbol{\varepsilon}$ and all further stochastic derivations are directly based on their definition. This stringent argumentation is regarded as a main advantage of this presentation.
@@
In total, this is an attempt for a general argumentation framework for model fitting with many possible applications and an introduction to the specific treatment for partially unpaired data. The focus of this presentation is on the applied researcher, explaining all derivations and results in detail as well as providing numerical implementation strategies and interpretations of numerical examples. Further work is encouraged in order to extend this framework or provide further examples of expressive applications.
6 Supplements
@@ Appendix A Completely Paired Data: Derivations for Errors in $\boldsymbol{y}$ only
Appendix B Completely Paired Data: Examples for Errors in $\boldsymbol{x}$ and $\boldsymbol{y}$
Derivation of Deming Regression
In order to derive the Deming regression, we need to solve
@@ Appendix B Completely Paired Data: Examples for Errors in $\boldsymbol{x}$ and $\boldsymbol{y}$
which one gets simply setting the gradient to zero
@@
which solution results in the classical Deming regression coefficients
@@
Fitting a Hyperplane and Gaussian Disturbance (Errors-In-Variables Multiple Linear Regression)
Further extending this argumentation to hyperplanes for $\boldsymbol{x}_{l}\in\mathbb{R}^{k}$ and $\boldsymbol{y}_{l}\in\mathbb{R}$, and ${\boldsymbol{\eta}}_{l}\sim\mathcal{N}(\boldsymbol{0},\,\sigma_{{\eta}}^{2}\cdot I_{{k}\times{k}})(\boldsymbol{s})\;\forall l=1,..,L$ for fitting an affine hyperplane model $M(\boldsymbol{x};\boldsymbol{\alpha})=\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{n}$ leads (utilizing Equation (4)) to the general optimization problem
[…]
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\prod\limits_{l=1}^{L}\;\text{e}^{-\frac{\left(\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{l,n}-y_{l}\right)^{2}}{2\,\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\right)\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}\right)}}$
=$\quad$= $\text{argmin}_{\;\boldsymbol{\alpha}}\quad\sum\limits_{l=1}^{L}\;\frac{\left(\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{l,n}-y_{l}\right)^{2}}{2\,\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\right)\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}\right)}\;.$
Appendix C Completely Unpaired Data: Examples for Errors in $\boldsymbol{x}$ and $\boldsymbol{y}$
The following examples are only the short summary of previously introduced examples translated to completely unpaired data:
Example: Fitting a Line and Gaussian Disturbance (Deming Regression)
Gaussian disturbances ${\eta}_{h}\sim\mathcal{N}(0,\,\sigma_{{\eta}}^{2})(s)\;\forall h=1,..,H$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$. Utilizing the one-dimensional affine model $M(x;\boldsymbol{\alpha})=\alpha_{1}+\alpha_{2}\cdot x$ and inserting it, we get
[…]
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\sum\limits_{l=1}^{L}\sum\limits_{h=1}^{H}\;\text{e}^{-\frac{(\alpha_{1}+\alpha_{2}\cdot x_{h}-y_{l})^{2}}{2\,(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2})}}$ (30)
Example: Fitting a Hyperplane and Gaussian Disturbance (Errors-In-Variables Multiple Linear Regression)
Gaussian disturbances ${\boldsymbol{\eta}}_{h}\sim\mathcal{N}(\boldsymbol{0},\,\sigma_{{\eta}}^{2}\cdot I_{k\times k})(\boldsymbol{s})\;\forall h=1,..,H$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$. Utilizing the $k$-dimensional affine model $M(\boldsymbol{x};\boldsymbol{\alpha})=\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{n}$ and inserting it, we get
[…]
=$\quad$= $\text{argmax}_{\;\boldsymbol{\alpha}}\quad\sum\limits_{l=1}^{L}\sum\limits_{h=1}^{H}\;\text{e}^{-\frac{\left(\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{h,n}-y_{l}\right)^{2}}{2\,\left(\left(\sum\limits_{n=1}^{k}\alpha_{n+1}^{2}\right)\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}\right)}}$ (32)
Example: Linear Interval Data Regression
Interval data given with
$\boldsymbol{x}_{h}$ $=\frac{1}{2}\left(\underline{\boldsymbol{x}_{h}}+\overline{\boldsymbol{x}_{h}}\right)$ (33)
$\boldsymbol{y}_{l}$ $=\frac{1}{2}\left(\underline{\boldsymbol{y}_{l}}+\overline{\boldsymbol{y}_{l}}\right)\;,$ (34)
and with $v_{h,i}:=\frac{1}{2}\left(\overline{x_{h,i}}-\underline{x_{h,i}}\right)$ and $w_{l,j}:=\frac{1}{2}\left(\overline{y_{l,j}}-\underline{y_{l,j}}\right)$
$\boldsymbol{\eta}_{h}$ $\sim\prod\limits_{i=1}^{k}U_{\left[-v_{h,i},v_{h,i}\right]}(s_{i})$ (35)
$\boldsymbol{\varepsilon}_{l}$ $\sim\prod\limits_{j=1}^{m}U_{\left[-w_{l,j},w_{l,j}\right]}(s_{j})\;.$ (36)
Utilizing the one-dimensional affine model $M(x;\boldsymbol{\alpha})=\alpha_{1}+\alpha_{2}\cdot x$ and inserting it, we get
$\Rightarrow$ $\;\text{argmax}_{\;\boldsymbol{\alpha}}\quad\sum\limits_{l=1}^{L}\sum\limits_{h=1}^{H}\;\frac{1}{v_{h}\cdot w_{l}}\cdot\text{max}\left[\;\text{min}\left[v_{h},c_{l,h,\text{max}}(\boldsymbol{\alpha})\right]-\text{max}\left[-v_{h},c_{l,h,\text{min}}(\boldsymbol{\alpha})\right],0\;\right]\;,$ (37)
with $c_{l,h,\pm}(\boldsymbol{\alpha}):=\frac{1}{2}\left(\underline{x_{h}}+\overline{x_{h}}\right)+\frac{1}{\alpha_{2}}\left(\alpha_{1}-\frac{1}{2}\left(\underline{y_{l}}+\overline{{y}_{l}}\right)\pm w_{l}\right)$ and $c_{l,h,\text{min}}=\text{min}(c_{l,h,\pm})$ and $c_{l,h,\text{max}}=\text{max}(c_{l,h,\pm})$.
Appendix D Numerical Implementation of the General Formula for Partially Unpaired Data
which increases computational efficiency significantly.
@@ Appendix D Numerical Implementation of the General Formula for Partially Unpaired Data
Appendix E Demonstration for a Plane Fit with Gaussian Disturbance for Partially Unpaired Data
@@
One possible part of demonstrating the flexibility of this framework is to show how it works in higher dimensions, which will be indicated by the previously introduced plane fit model. Using $L=H=1600$ data points $\boldsymbol{x}_{h}\in\mathbb{R}^{2}$ and $y_{l}\in\mathbb{R}$ and numbers of subgroups $R\in\{6,18,36,L=H\}$. The data generation parameters are the same as for the base scenario $A$ in the line fit section but with the correct values $\alpha_{1}=0$, $\alpha_{2}=0.2$ and $\alpha_{3}=0.4$. In Figure 7 the fitting results are presented.
@@
Obviously the partitioning of the data has much more possibilities due to a much richer neighboring information for $\boldsymbol{x}_{h}$ in 2D. In consequence, the discussed dissimilarity of the groups to the total data set gets more difficult to study. Nonetheless, it can be demonstrated that the utilization with more pairing information does increase accuracy but that practically there might be a saturation for this.
@@
Figure 7: Illustration of example results for different plane fitting scenarios. Four different scenarios of partial pairing with $R=6$, $R=18$, $R=36$ and $R=L=H$ unpaired subgroups. For each case: Top plot: representing the unpaired data subgroups by color-coding. Bottom plot: presenting the fitted plane (red) in the full data (gray).
Appendix F Application of $R_{\delta}^{2}$

### v2

@@ 2.4.1 Example: Fitting a Line and Gaussian Disturbance
which represents a solution to the Deming type problem for partially unpaired data.
@@
Gaussian disturbances in input and output variables with
${\boldsymbol{\eta}}_{h}\sim\mathcal{N}(\boldsymbol{0},\,\text{diag}(\sigma_{{\eta},1}^{2},..,\sigma_{{\eta},k}^{2}))(\boldsymbol{s})\;\forall h=1,..,H$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$, and utilizing the $k$-dimensional affine model $M(\boldsymbol{x};\boldsymbol{\alpha})=\alpha_{1}+\sum\limits_{n=1}^{k}\alpha_{n+1}\cdot x_{n}$ and inserting it, we get by applying Equation (11)
@@ 2.4.3 Example: Linear Interval Data Regression
As final example, we present interval data that are given with $\boldsymbol{x}_{h}:=\frac{1}{2}\left(\underline{\boldsymbol{x}_{h}}+\overline{\boldsymbol{x}_{h}}\right)$ and $\boldsymbol{y}_{l}:=\frac{1}{2}\left(\underline{\boldsymbol{y}_{l}}+\overline{\boldsymbol{y}_{l}}\right)$, and $v_{h,i}:=\frac{1}{2}\left(\overline{x_{h,i}}-\underline{x_{h,i}}\right)$ and $w_{l,j}:=\frac{1}{2}\left(\overline{y_{l,j}}-\underline{y_{l,j}}\right)$, and
@@ 2.5 Extensions
2.5.1 Numerical Implementation of the General Formula for Partially Unpaired Data
which increases computational efficiency significantly.
@@ 3.1 Demonstration for a Line Fit
See Figure 2 for presentation of example objective function and resulting line fits of scenario $A$ and $D$.
@@
The columns correspond to different numbers of unpaired subgroups $R\in\{1,3,12,L\}$. For the completeley unpaired case ($R=1$) only the objective function is presented.
@@
In Supplement C a plane fit is presented as an example for 2D input variables.
@@ 3.2 Nonlinear Model with Anisotropic Observation Errors
is presented with anisotropic Gaussian disturbances $\sigma_{\eta}=0.2$ and $\sigma_{\varepsilon}=0.1$ for $L=H=300$ data points. In this model, nonlinearity holds with respect to $x$ and not with respect to $\boldsymbol{\alpha}$, which is deliberate and not necessary. In Figure 4 the results are presented for a completely paired case ($R=L=H$) (left) and a partially unpaired case (right) with two areas of lost pairing information. The general algorithm was implement according to section 2.5.1 with the simple trapezoidal rule for integration and the Nelder-Mead-optimization in Matlab.
@@ 4 Real Data Study: Life Expectancy
For evaluation purposes, we perform a train-test-split with $172$ training countries and $20$ test countries in order to judge if we can learn from the training countries the life expectancy for the test countries based on the input variables. A major challenge is that the goodness of fit cannot be measured with the classical $R^{2}$ since it is only valid for cases with no errors in the input variables. In consequence, we utilize the extension of $R^{2}$ to the errors-in-variables approach $R_{\delta}^{2}$ [Cheng et al., 2014] for multiple linear regression (see Supplement D for more details).
First, we consider the case of the model fit with perfect pairing ($R=L=H$). Since the proposed algorithm was implemented with Monte Carlo methods as described in Section 2.5.1, it was verified with an explicit solution for the multiple linear regression model [Cheng et al., 2014] and showed equivalent results. The results are presented in Figure 6A as correlation and error plots of predicted and real output values for the train and test data sets. They show that the training allows a high quality prediction for the test countries with the multiple linear regression model for life expectancy.
@@
Second, one major approach in this work is to investigate explanatory power of the model fit if the train data is partially unpaired, i.e. if we only consider data of groups of countries rather than individual countries for model training. To achieve this, we introduce the country grouping along an additional criteria which is not part of the input: GDP per capita (current USD) [NY.GDP.PCAP.CD] with increasing GDP per capita for each group. Please note, although the grouping is performed by GDP, the GDP itself is not part of the predictors and the overall information level for the predictors is reduced by this grouping compared to the completely paired data set. We considered $3$ different groupings with group sizes $L_{r}=H_{r}$ (the last group always consists of residual countries) of $4$ ($R=44$), $8$ ($R=22$) and $16$ ($R=11$). The results of the model fit based on this country groups are presented in Table 1.
@@
The $R_{\delta}^{2}$ for train and test data is presented which show overall very high goodness-of-fit values but with decreasing values for increasing group sizes, as one might expect. Please note, the $R_{\delta}^{2}$ for testing contains a rather strong noise component since it is calculated by only $20$ data points in the test set. In addition, in Figure 6B (for group size $16$) it is demonstrated that the fit results still align well for training and testing even for $11$ subgroups. It is noted, that the grouping is part of the training to get $\boldsymbol{\alpha}$, and not part of these scatter plots. This indicates that the model fit and the prediction for the test countries work quite well based only on the country groups, confirming that the pairing information can be reduced and this still provides a valuable model fit.
@@ 5 Discussion and Conclusion
Specifically, we regard the framework presented as a generalization of total least squares approaches [Markovsky and Van Huffel, 2007] as a more general errors-in-variables treatment, since we can reproduce classical results of total least squares by utilizing Gaussian errors and a line model. On the other hand, we are not restricted to these assumptions, e.g. allowing for interval data using uniform distributions. We want to emphasis, that the only initial random variables of this paper are the uncertainty random variables $\boldsymbol{\eta}$ and $\boldsymbol{\varepsilon}$ and all further stochastic derivations are directly based on their definition. This stringent argumentation is regarded as a main advantage of this presentation.
@@
In total, this is a general argumentation framework for model fitting with many possible applications and an introduction to the specific treatment for partially unpaired data. The focus of this presentation is on the applied researcher, explaining all derivations and results in detail as well as providing numerical implementation strategies and interpretations of numerical examples. Further work is encouraged in order to extend this framework or provide further examples of expressive applications.
@@ Appendix A Completely Paired Data: Derivations for Errors in $\boldsymbol{y}$ only
Appendix B Completely Paired Data: Relation to Deming Regression
Deming regression is equivalent to the maximum likelihood for independent normally distributed observation errors in $x_{l}$ and $y_{l}$ ($l=1,..,L$), i.e. ${\eta}_{l}\sim\mathcal{N}(0,\,\sigma_{{\eta}}^{2})(s)\;\forall l=1,..,L$ and ${\varepsilon}_{l}\sim\mathcal{N}(0,\,\sigma_{{\varepsilon}}^{2})(s)\;\forall l=1,..,L$ with a line model. Since the true value $s_{l}$ for $x_{l}$ is unknown just like the parameters $\boldsymbol{\alpha}$ they are estimated by maximizing them simultaneously with the parameters $\boldsymbol{\alpha}$, which leads to the effect that the estimation of the true values influences the estimation of the parameters $\boldsymbol{\alpha}$:
$\text{argmax}_{\;\boldsymbol{\alpha},s_{1},..,s_{L}}\quad\prod\limits_{l=1}^{L}\;\text{e}^{-\frac{1}{2\,\sigma_{{\varepsilon}}^{2}}(y_{l}-\alpha_{1}-\alpha_{2}\cdot s_{l})^{2}-\frac{1}{2\,\sigma_{{\eta}}^{2}}(x_{l}-s_{l})^{2}}$
In the case of this paper, we are also estimating the best parameters $\boldsymbol{\alpha}$ but independently of any specific true value $s_{l}$: We are averaging over all possible true values by the use of the law of total probability, compare Equation (5), which can be interpreted as an integrated Deming regression. This is a valid alternative perspective leading to a slightly more defensive estimation of $\boldsymbol{\alpha}$ which is not influenced by the estimation of $s_{l}$. An interesting observation is that the second part of derived objective function in Equation (6)
@@ Appendix B Completely Paired Data: Relation to Deming Regression
actually leads to the classical Deming equations and the first part of Equation (6) $\frac{L}{2}\,\ln\left(\alpha_{2}^{2}\,\sigma_{{\eta}}^{2}+\sigma_{{\varepsilon}}^{2}\right)$ can be interpreted as a penalty added to the classical Deming regression, showing the more defensive estimation approach. Especially for large $\sigma_{\eta}$ the parameter $\alpha_{2}$ will tend slightly more to zero. The derivation of the classical Deming regression in this context is directly the minimization of Equation (27) by setting the gradient to zero
@@
whose solution results in the classical Deming regression coefficients
@@
In conclusion, the presented approach leads for Deming type problems to formulas, which we call integrated Deming regression. These formulas can be interpreted as penalized classical Deming regression showing the more defensive approach by averaging over the true value during estimation of $\boldsymbol{\alpha}$ compared to estimating them simultaneously in classical Deming regression.
@@
Appendix C Demonstration for a Plane Fit with Gaussian Disturbance for Partially Unpaired Data
@@
One possible part of demonstrating the flexibility of this framework is to show how it works in higher dimensions, which will be indicated by the previously introduced plane fit model. We are using $L=H=1600$ data points $\boldsymbol{x}_{h}\in\mathbb{R}^{2}$ and $y_{l}\in\mathbb{R}$ and numbers of subgroups are $R\in\{6,18,36,L=H\}$. The data generation parameters are the same as for the base scenario $A$ in the line fit section but with the true values $\alpha_{1}=0$, $\alpha_{2}=0.2$ and $\alpha_{3}=0.4$. In Figure 7 the fitting results are presented.
@@
Obviously the partitioning of the data has much more possibilities due to a much richer neighboring information for $\boldsymbol{x}_{h}$ in 2D. In consequence, the discussed dissimilarity of the groups to the total data set gets more difficult to study. Nonetheless, it can be demonstrated that the utilization with more pairing information does increase accuracy and meaningful estimation can be performed even with a low number of groups.
@@
Figure 7: Illustration of example results for different plane fitting scenarios. Four different scenarios of partial pairing with $R=6$, $R=18$, $R=36$ and $R=L=H$ unpaired subgroups. For each case: Left plot: representing the unpaired data subgroups by color-coding. Right plot: presenting the fitted plane (red) in the full data (gray).
Appendix D Application of $R_{\delta}^{2}$